### v2

# Dynamical constraints on the vertical structure of Jupiter’s polar cyclones

Abstract
Jupiter’s poles feature striking polygons of cyclones that drift westward over time, a motion governed by $\beta$-drift (vortex motion caused by the latitudinal variation of the Coriolis force). This study investigates how $\beta$-drift and the resulting westward motion depend on the depth of these cyclones. Counterintuitively, shallower cyclones drift more slowly, a consequence of stronger vortex stretching. By employing a 2D quasi-geostrophic model of Jupiter’s polar regions, we constrain the cyclones’ deformation radius, a key parameter that serves as a proxy for their vertical extent, required to replicate the observed westward drift. We then explore possible vertical structures and the static stability of the poles by solving the eigenvalue problem that links the 2D model to a 3D framework, matching the constrained deformation radius. These findings provide a foundation for interpreting upcoming Juno microwave measurements of Jupiter’s north pole, offering insights into the static stability and vertical structure of the polar cyclones. Thus, by leveraging long-term motion as a novel constraint on vertical dynamics, this work sets the stage for advancing our understanding of the formation and evolution of Jupiter’s enigmatic polar cyclones.
1Department of Earth and Planetary Sciences, Weizmann Institute of Science, Rehovot, Israel
nimrod.gavriel@weizmann.ac.il
Preprint November 2, 2025
This article has been published in
Proceedings of the National Academy of Sciences (PNAS) (2025), 122 (44) e2503737122.
The final version is available at DOI:10.1073/pnas.2503737122.
(Received Feb 19, 2025; Revised June 24, 2025; Accepted September 16, 2025)
Key Points
  • $\beta$-drift of cyclones, linked to Jupiter’s polar cyclones’ westward drift, strengthens for deeper cyclones as vortex stretching diminishes.
  • A 2D quasi-geostrophic model estimates a deformation radius of about 220 km for the north pole and 360 km for the south pole.
  • The estimated deformation radii uncover possible vertical structures shaped by vertical stability.
Plain language summary
Jupiter’s poles feature clusters of large, long-lasting cyclones that slowly drift westward. We argue that the depth to which these cyclones extend below the cloud layers significantly influences this drift rate. By comparing computer simulations with observations, we estimate the depth of the cyclones and shed light on the stability of Jupiter’s atmosphere at high latitudes. These insights are particularly critical for Juno’s extended mission, which now allows the Microwave Radiometer (MWR) to observe these cyclones with unprecedented resolution, providing a direct test of our framework and facilitating quantitative interpretation of their vertical structure. Our results directly support this new phase of the Juno mission, offering key constraints to understand the formation, persistence, and dynamics of these giant polar storms.
1 Introduction
Figure 1: Impact of Stretching on $\beta$-Drift Dynamics in a Single-Layer QG Framework: A Single Cyclone on a $\beta$-Plane. (a–b) Evolution of the zonal ($u_{\beta}$) and meridional ($v_{\beta}$) drift velocities over time for four different values of the Burger number and $b=1.2$. Dashed curves show exponential fits with amplitude $\rm{U}_{\beta}$ and timescale $\tau$. Here, $(u_{\beta},v_{\beta})$ are non-dimensional, scaled by $\hat{\beta}V$. (c) Plots of $\rm{U}_{\beta}$ (solid line, left axis) and $\tau$ (dashed line, right axis, in rotation periods) as functions of $\rm Bu$} for three values of the shape factor $b$ (legend in panel d). The exponential parameterization becomes inadequate for ${\rm Bu}\gtrsim 5$ (see Fig. S2). (d) Solid line: Phase angle $\alpha$ (defined in panel e) between $u_{\beta}$ and $v_{\beta}$ as a function of $\rm Bu$}. Dashed line: Strength of the $\beta$-gyres, quantified as $\Delta\psi_{g}=\max(\psi_{g})-\min(\psi_{g})$, with values non-dimensionalized by $\hat{\beta}RV$. (e) The vorticity field ($\xi_{\rm g}$, in color), streamfunction ($\psi_{\rm g}$, in contours), and velocity field ($\mathbf{u}_{\rm g}$, arrows) after five rotation periods for $\rm{Bu}=0.59$ and $b=1.2$. The $\beta$-gyres appear as red and blue lobes surrounding the cyclone core. The colorbar represents the scaled (by $\hat{\beta}V/R$) vorticity field. See Movie S1 for the simulation.
Probing beneath the clouds of Jupiter
Historically, our understanding of Jupiter’s atmosphere has been limited to observations of its visible cloud layer, located near the 1-bar pressure level. This limitation arises from the challenges of probing beneath the optically thick ammonia and water clouds Atreya et al. (1999); de Pater et al. (2016). Until recently, the only direct measurements of Jupiter’s subcloud layers were the localized and costly in-situ data collected by the Galileo Probe Young (1998). However, an indirect method using gravity measurements (by tracking changes in the Juno spacecraft’s acceleration during close flybys) has enabled insights into the large-scale flows at depth. This approach has revealed the 3D structure of zonal flows within Jupiter’s subequatorial jet streams Kaspi et al. (2018); Galanti et al. (2021); Kaspi et al. (2023) and constrained the depth of the Great Red Spot (GRS) to less than $500$ km Galanti et al. (2019); Parisi et al. (2021).
Another approach to exploring Jupiter’s subcloud layers involves observations at cloud-penetrating frequencies. For example, using radio frequencies, the Earth-based Very Large Array (VLA) telescopes have probed down to pressures of approximately 8 bars de Pater et al. (2016). The Juno Microwave Radiometer (MWR) offers an enhanced use of cloud-penetrating frequencies, leveraging its orbital proximity to achieve deeper measurements with finer spatial resolution. Equipped with six channels, the MWR has been providing spatially resolved observations of brightness temperature down to depths of approximately 240 bars Bolton et al. (2017); Janssen et al. (2017). These thermodynamic measurements are influenced both by temperature variations Fletcher et al. (2021) and the opacity of ammonia clouds Li et al. (2017). Although there is ongoing debate about the relative contributions of these factors, the observed deep patterns reveal valuable information about the vertical structure and depth of atmospheric phenomena seen at the cloud tops. For example, the MWR detected alternating zonal bands of ammonia depletion and enrichment extending to 240 bars between midlatitude jets. These patterns correspond to the upwelling and downwelling branches expected in a Ferrel-cell-like circulation Duer et al. (2021, 2023).
Another key instrument on Juno, the Jovian Infrared Auroral Mapper (JIRAM), was designed to study the aurora and atmospheric chemistry at 5-7 bar Adriani et al. (2017). During Juno’s polar orbits, as JIRAM is insensitive to half the pole being dark and can wholly observe the pole, it discovered a new phenomenon: a polar "crystal" of massive cyclones (each $\sim 5,000$ km wide) inhabiting each pole, consisting of a polar cyclone (PC) and 8 circumpolar cyclones (CPCs) at the north pole, and 5 at the south Adriani et al. (2018). The MWR instrument, designed to function optimally at the shorter distances from Jupiter’s cloud-tops during the around-equatorial skims in its original polar orbit, now enters a phase in Juno’s extended mission, where the orbit has considerably shifted, allowing MWR measurements in near north pole of Jupiter Orton et al. (2024). This enables multi-channel measurements with sufficient resolution to capture the north polar cyclones, similar to analogous previous measurements of vortices in the midlatitudes Bolton et al. (2021). This study aims to provide a theoretical framework for interpreting the upcoming MWR observations of Jupiter’s north pole, offering predictions about the vertical structure and extent of the polar cyclones.
Jupiter’s Polar Cyclones
The energetic processes sustaining Jupiter’s polar cyclone configurations, while explored in prior studies O’Neill et al. (2015); Brueshaber et al. (2019); Hyder et al. (2022); Siegelman et al. (2022b, a), remain largely unconstrained. Nevertheless, substantial progress has been made in understanding the principal momentum balance that governs these cyclones, explaining both their stability and long-term motion. This momentum balance is closely related to the concept of $\beta$-drift, a secondary effect that causes vortices to move due to their interaction with a background vorticity gradient Sutyrin and Flierl (1994) known to influence the motion of tropical cyclones on Earth Franklin et al. (1996). The phenomenon of $\beta$-drift has been explored extensively in idealized settings, including theoretical studies Adem (1956); Smith and Ulrich (1990), numerical simulations Lam and Dritschel (2001); Gavriel and Kaspi (2023), and laboratory experiments Firing and Beardsley (1976); Carnevale et al. (1991); Stegner and Zeitlin (1998); Flór and Eames (2002); Benzeggouta et al. (2025).
Traditionally, as suggested by the term $\beta$-drift", studies have focused on the interaction of cyclones with $\beta$, the meridional gradient of the planetary vorticity $f$, causing a poleward-westward migration through the generation of opposite vorticity anomalies (usually termed $\beta$-gyres) surrounding the core. However, in addition to $\beta$, cyclones on Jupiter’s poles exhibit mutual repulsion due to the interaction of the vorticity gradient of one cyclone with the velocity field of another. This mutual interaction allows for the stable crystal-like configuration of the cyclones Gavriel and Kaspi (2021), provided that they are surrounded by an anticyclonic annulus, commonly referred to as shielding Li et al. (2020). This repulsion also explains the oscillatory motion of the polar cyclones Gavriel and Kaspi (2022), as observed by Juno Mura et al. (2021).
Beyond oscillatory motion, the cyclones exhibit a mean westward drift, with rates of approximately $3^{\circ}$ and $7^{\circ}$ per year at the north and south poles, respectively Adriani et al. (2020); Mura et al. (2022). This westward drift arises from the $\beta$-drift acting collectively on the group of cyclones. By adopting a "center of mass" perspective, the mutual interactions between the cyclones are effectively averaged out, leaving only the cumulative interaction of the group with $\beta$. This dynamic leads to a group oscillation around the poles accompanied by a westward precession, manifesting as the observed drift Gavriel and Kaspi (2023).
In this study, we investigate how the vertical structure and depth of Jupiter’s polar cyclones influence their mean westward drift. First, we employ an idealized 2D model to explore how depth, and the associated column stretching, affects $\beta$-drift on a vortex in a $\beta$-plane. Next, we perform a parameter scan using a 2D Quasi-Geostrophic (QG) polar model with Jupiter’s polar cyclones initialized in the system to constrain the QG deformation radius ($L_{d}$), a proxy for depth, such that the modeled westward drift matches observational data. Finally, we solve the eigenvalue problem posed by comparing 2D and 3D QG systems to deduce the vertical structures and static stability of the polar cyclones based on the constrained $L_{d}$.
Figure 2: Constraining $L_{d}$ at Jupiter’s Poles Using Observed and Simulated Cyclone Drifts. (a,b) Observed trajectories of the north (a) and south (b) polar cyclones over five years Mura et al. (2022). The symbols $\odot$ and $\otimes$ mark the start and end points, respectively. The background is a JIRAM infrared image from PJ4 (Image from Adriani et al. (2018)), corresponding to the onset of these trajectories. (c,d) Simulated trajectories after five years for a model that reproduces the observed mean westward drift. The color scale indicates the magnitude of the velocity vector, while the streaks represent the flow direction through line integral convolution (LIC) of the velocity field at $t=0$. (e–j) Sample model trajectories from the nested sampling routine, illustrating diverse outcomes. The numbers above the panels are the mean and standard deviation (between the cyclones) of the westward drift. The parameters used in each panel are listed below the panels. (k–n) Histogram plots of the posterior parameter distributions produced by Ultranest Buchner (2021), after sampling 15,00015{,}000 parameter sets per pole for the north (blue) and south (orange) polar cyclones. The dashed lines denote the median, and the dotted lines mark $\pm 1\sigma$ for the distributions. These values are written above the panels. Markers represent the parameter values of the successful models (panels c,d). For a corner plot of the joint posterior distributions, see Fig. S4.
The Impact of Vertical Depth on Vortex $\beta$-Drift
A standard framework for understanding the motion of vortices due to $\beta$-drift is the barotropic QG equation of a seeded vortex on a $\beta$-plane Smith and Ulrich (1990); Sutyrin and Flierl (1994); Smith (1991). In this formulation, cyclones undergo poleward-westward propagation in a planetary $\beta$, with the drift amplitude scaling with the dimensionless parameter $\hat{\beta}=\frac{\beta R^{2}}{V}$ Smith (1997); Gavriel and Kaspi (2023), where $R$ and $V$ are the characteristic size and velocity of the cyclone. In this section, we seek to understand how the vertical extent (through stretching) modifies the classical $\beta$-drift by adding the stretching term to the QG equation, before treating other important complexities such as vortex-vortex interactions. Stretching, in this context, refers to the vertical elongation (or compression) of a fluid column accompanied by horizontal convergence (or divergence). In QG dynamics, such vertical motions alter the potential vorticity (PV) by concentrating or diluting the background vorticity (see Fig. S1 and the accompanying text for more intuition).
Following the derivation in Smith and Ulrich (1990) but including stretching (see SI for scaling, derivation, and other details), we develop the equations in a reference frame moving with the center of the cyclone. The scaled resulting QG equation becomes (focusing on the dominant terms)
$\frac{\partial}{\partial t}\left(\xi_{\rm g}-\frac{\psi_{\rm g}}{\rm Bu}\right)=-\hat{\beta}v_{\rm v}-\mathbf{u}_{\rm v}\cdot\mathbf{\nabla}\xi_{\rm g},$ (1)
where $\xi$ and $\psi$ are the vorticity and streamfunction, respectively, and $\mathbf{u}=(u,v)$ is the velocity vector. Subscripts $\rm v$ and $\rm g$ represent, respectively, a constant background vortex flow and a "generated" field leading to the $\beta$-drift. The Burger number, ${\rm Bu}\equiv\left(\frac{L_{d}}{R}\right)^{2}$, incorporates the effect of the deformation radius, $L_{d}=\frac{\sqrt{gH}}{f_{0}}$, where $g$ is gravitational acceleration, $H$ is the layer depth, and $f_{0}$ is the Coriolis parameter. The term $\frac{\partial}{\partial t}\left(\frac{\psi_{\rm g}}{\rm Bu}\right)$ captures the influence of stretching in the single-layer QG model.
The model is solved for vortex profiles of the dimensional form Chan and Williams (1987):
$\xi_{\rm v}=\frac{V}{R}\left(2-\left(\frac{r_{\rm c}}{R}\right)^{b}\right)\exp\left[\frac{1}{b}\left(1-\left(\frac{r_{\rm c}}{R}\right)^{b}\right)\right],$ (2)
where $r_{\rm c}$ is the distance from the vortex’s center, and $b$ is a shape factor that controls the sharpness and shielding of the vortex profile. For a more detailed analysis of the role of $b$, see Li et al. (2020). The initial-value problem is solved using the Dedalus PDE solver Burns et al. (2020). The results from this model, as a function of $\rm Bu$}, are shown in Fig. 1. The zonal $\beta$-drift velocity, defined as $u_{\beta}\equiv u_{\rm g}{\mid_{(x=0,y=0)}}$, displays an exponential relaxation with time, which is modeled by $\rm{U}_{\beta}\bigl(1-e^{-t/\tau}\bigr)$ (Fig. 1a), a parameterization that provides a good fit for $\rm{Bu}\lesssim 5$.
We further examine the dependence of the fitted maximum drift velocity ($\rm{U}_{\beta}$) and timescale ($\tau$) on layer depth (Fig. 1c) for three values of the shape factor $b$ within a physically relevant range. The results indicate that while the drift amplitude is relatively insensitive to changes in $b$, it becomes more negative (more strongly westward) as the Burger number increases (or equivalently, as $H$ increases).
In Gavriel and Kaspi, 2023 Gavriel and Kaspi (2023) (their Fig. 2), the layer depth $H$ is assumed to be infinite, which implies $L_{d}\rightarrow\infty$ and $\rm Bu$$\rightarrow\infty$, effectively removing the stretching term from Eq. 1. This assumption highlights that for very deep layers, any change in the fluid upper surface height $(\delta h)$ is negligible compared to the total depth $(\frac{\delta h}{H}\ll 1)$, rendering stretching effects unimportant to the dynamics.
However, when $H$ (and thus $\rm Bu$}) is finite, the PV anomaly generated by the $\beta$-effect (the term $\hat{\beta}v_{\rm{v}}$) is divided between two parts: the formation of $\beta$-gyres (through $\frac{\partial\xi_{\rm g}}{\partial t}$) and column stretching (through $\frac{\partial}{\partial t}\frac{\psi_{\rm g}}{\rm Bu}$). As stretching has more relative significance in shallower layers, it claims a larger share of the total PV response, thereby weakening the $\beta$-gyres. This trend is reflected in Fig. 1c, where shallower layers correlate with stronger stretching and hence smaller $\beta$-drift velocities. This reduction is also evident in Fig. 1d, where the difference between the extreme streamfunction values of the $\beta$-gyres, $\Delta\psi_{g}$, becomes smaller.
Using the Jovian values for the north and south poles of $\hat{\beta}$, $R$, and $V$ from Gavriel and Kaspi (2023), we find that matching the observed westward drift corresponds to $\rm{U}_{\beta}=-0.46$ for the north pole and $\rm{U}_{\beta}=-0.54$ for the south pole, marked in Fig. 1c. These values correspond to ${\rm Bu}=0.49$ for the north pole and ${\rm Bu}=0.59$ for the south, which is used in Fig. 1e to show the generated fields after five rotation periods. The influence of $\rm Bu$} extends beyond the reduction of the $\beta$-drift amplitude. Smaller $\rm Bu$} values also reduce the time required to reach the peak drift velocity ($\tau$ in Fig. 1c) and lead to a more zonally directed drift overall ($\alpha$ in Fig. 1d).
From ${\rm Bu}=(0.49,0.59)$ and $R=(811,861)$ km Gavriel and Kaspi (2023), the resulting deformation radii are $L_{d}=(568,658)$ km for the north and south poles, respectively. These values should not be interpreted as definitive physical estimates, but rather as outcomes of an idealized model that omits vortex–vortex interactions and the polar variation of $f$. Instead, this simplified framework is meant to provide insight into the role of stretching alone in modulating $\beta$-drift. In the next section, we consider these additional complexities to derive more realistic values of $L_{d}$ needed to match the observed drift rates.
Estimating the Deformation Radius Using a Single-Layer QG Model of the Jovian Poles
While the previous section focused on how vertical extent modifies $\beta$-drift, the idealized model used there does not provide a realistic estimate of $L_{d}$. To address this limitation, we now present a single-layer QG model initialized with a polar vortex crystal. This model evolves under the full spherical form of the vertical component of Jupiter’s rotation rate, $f=2\Omega\cos\left(\frac{r}{R_{\rm J}}\right)$, where $\Omega$ is Jupiter’s rotation rate, $R_{\rm J}$ is Jupiter’s radius, and $r$ is the distance from the pole. We use Dedalus Burns et al. (2020) to solve the polar QG equation Vallis (2017):
$\frac{D}{Dt}\left(\nabla^{2}\psi+2\Omega\cos\left(\frac{r}{R_{\rm J}}\right)-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (3)
with multiple cyclones prescribed by Eq. 2, with one PC and eight (five) CPCs placed initially at the observed locations during Juno’s PJ4 orbit for the north (south) poles (Fig. 2a-b). For simplicity and because of limited observational data, we assume identical cyclones. Due to this choice and the absence of small-scale forcing, the oscillatory motion discussed by Gavriel and Kaspi (2022) is not excited. Hence, we focus on reproducing the mean westward drift. While zonal jets could, in principle, influence the drift, no such flows have been observed in the vicinity of the cyclones Rogers et al. (2022). Moreover, JunoCam imagery is consistent with a two-dimensional, isotropic turbulent regime Orton et al. (2017), and thus zonal flows are not included in this model.
This model involves four unknown parameters: $L_{d}$, and the three cyclone profile parameters $(R,V,b)$. To explore the phase space of this model, we use the Ultranest library Buchner (2021), which employs a Bayesian nested sampling approach to identify the best-fitting solutions Buchner (2023). As a benchmark, we use the 5-year observed trajectories of the north and south polar cyclones Mura et al. (2022) (Fig. 2a-b) to calculate the average westward drift between the cyclones. This value is then used to calculate a likelihood function that evaluates the probability of a parameter set being accurate, based on the resultant model westward drift (see SI for details on the numerical model, likelihood function, and model sampling).
We sampled 15,00015{,}000 parameter sets for each pole, integrating each simulation for seven years and calculating the resultant mean westward drift between the cyclones during the last five years (where the first 2 years act as a spin-up). Illustrative model outputs (Fig. 2 panels e–g for the north pole and panels h-j for the south) span a variety of outcomes: drifts too fast (panels e,h), too slow (panels f and i), eastward (panel g), and mergers (panel j). Examples of parameter sets that successfully recreate the observed drift are shown in Fig. 2c-d.
Due to intrinsic parameter degeneracies, the algorithm returns a posterior distribution rather than a specific set of parameters that matches the observed westward drift. From this distribution, we obtain a more realistic estimate for the deformation radius of $L_{d}=225.58_{-46.93}^{+83.94}$ km and $L_{d}=363.69_{-44.33}^{+74.50}$ km at the $1\sigma$ confidence level for the north and south poles (Fig. 2k), respectively. The resulting posterior distributions for $R$, $V$, and $b$ (Fig. 2l–n) align well with plausible mean values for Jupiter’s polar cyclones, as estimated from cloud feature tracking at the poles Grassi et al. (2018).
Figure 3: Interpretation of the Cyclones’ Vertical Structure from the Deformation Radius. The first row corresponds to the "mode 0" solutions of the eigenvalue problem described by Eq. 6, while the second row corresponds to "mode 1". (a,d) Vertical profiles of $\Phi_{n}$ for each mode. (b,e) Solutions for $L_{d}$ (green shades) for various combinations of NN (Brunt-Väisälä frequency) and $H$ (depth). The estimated $L_{d}$ distributions from Fig. 2k are shown by the blue and orange curves for the north and south poles, respectively. (c,f) Solutions for $\Phi_{n}$ as a function of pressure below the cloud layer, following the NN and $H$ relationship indicated by the blue solid curve in panels (b) and (e). Green horizontal dotted lines indicate the depths at which each MWR channel has maximum sensitivity (Janssen et al., 2017), though each channel is sensitive over a broader range of depths. The numbers on the right ordinate show the altitude relative to 1 bar. Grey dashed lines in panels (b,c,e,f) denote the values used to generate the profiles in panels (a) and (d), and to produce Fig. 4. The color along the grey dashed line in panels (c,f) corresponds to the $\Phi(z)$ profiles shown in panels (a,d). The purple dashed line in panel (c) represents the theoretical scaling $H\sim\mathrm{Ro}^{1/2}Rf_{0}/N$ (adapted from (Aubert et al., 2012)), using representative values of $R=800\,\rm{km}$ and $U=90\,\rm{ms}^{-1}$ from Fig. 2m-n.
Interpreting the Vertical Structure of Jupiter’s Polar Atmosphere from the Drift Rate of the Polar Cyclones
The standard shallow-water (SW) expression for the deformation radius, $L_{d}=\sqrt{gH}/f_{0}$, is not well-suited to Jupiter’s upper atmosphere. This expression assumes a constant density and no vertical dependence of the flow, whereas Jupiter’s upper-atmospheric density increases exponentially with depth Guillot and Morel (1995) over many scale heights. To account for these vertical variations, we consider the continuously stratified QG vorticity equation Vallis (2017):
$\frac{D}{Dt}\left(\nabla^{2}\psi+f+\frac{f_{0}^{2}}{\tilde{\rho}}\frac{\partial}{\partial z}\left(\frac{\tilde{\rho}}{N^{2}}\frac{\partial\psi}{\partial z}\right)\right)=0,$ (4)
where $N$ is the Brunt-Väisälä frequency, indicating the stability of the atmosphere with respect to vertical motion, and $\tilde{\rho}$ is the vertical density profile.
To account for variations in the vertical direction, we assume the streamfunction can be separated into horizontal and vertical components:
$\psi(x,y,z,t)=\sum_{n}\psi_{n}(x,y,t)\Phi_{n}(z)$ (5)
where $\Phi_{n}(z)$ is the vertical profile for each $n$. Substituting this form into the QG equations requires each $\Phi_{n}$ to satisfy the vertical eigenvalue problem Smith and Vallis (2001):
$\frac{f_{0}^{2}}{\tilde{\rho}}\frac{\partial}{\partial z}\left(\frac{\tilde{\rho}}{N^{2}}\frac{\partial\Phi_{n}}{\partial z}\right)+\Gamma_{n}\Phi_{n}=0,$ (6)
where $\Gamma_{n}$ is an eigenvalue. For each mode $n$, the corresponding deformation radius is given by $L_{d,(n)}=\Gamma_{n}^{-\frac{1}{2}}$ (see SI for the full derivation). Assuming that the cyclones are dominated by a single vertical mode, the horizontal dynamics derived from the two-dimensional QG equation (Eq. 3), and the three-dimensional QG equation (Eq. 4) will yield similar results for the given $L_{d,(n)}$; the linearized dynamics are identical (see SI). Given our estimates of the $L_{d}$ values required to achieve the observed westward drift (Fig. 2), we can use the eigenvalue problem posed by Eq. 6 to investigate which vertical structures (eigenfunctions $\Phi_{n}(z)$) satisfy the required $L_{d}$.
We solve this eigenvalue problem using the Dedalus library Burns et al. (2020), applying boundary conditions of no vertical velocity at the tropopause $(\partial_{z}\Phi_{n}|_{{}_{z=0}}=0)$ and no flow at the bottom $(\Phi_{n}|_{{}_{z=-H}}=0)$. For $\tilde{\rho}(z)$, we adopt a reference density profile derived from interior models of Jupiter, constrained by a neural-network-based parameter sweep Ziv et al. (2024a, b). We focus here on the two lowest modes ($n=\{0,1\}$), which are illustrated in Fig. 3(a,d). The $n=0$ mode corresponds to a simple decay profile, whereas $n=1$ changes sign along the column. This focus on the two leading modes aligns with estimates from Earth’s oceans, where they capture the bulk of the kinetic energy Smith and Vallis (2001); Ferrari and Wunsch (2010); higher modes require excitations with elaborate vertical profiles.
Since Eq. 6 depends only on $H$ and $N$ (here assumed constant, see discussion in the SI), we present $L_{d,(n)}$ solutions as a function of these two parameters in Fig. 3(b,e), for both modes. The constrained range of deformation radii from the previous section (Fig. 2k) is indicated for the north (blue) and south (orange) poles, essentially giving an $H$ as a function of $N$ for each pole. Focusing on the relationship defined by the blue solid curve in Fig. 3(b,e), we plot the corresponding $\Phi_{n}(z)$ profiles as a function of pressure below the cloud level for different $N$ values (Fig. 3(c,f)).
Interestingly, in agreement with the eigenvalue results, the $H$ vs $N$ line in Fig. 3c closely follows a scaling law derived for the aspect ratio of vortices in stratified fluids, supported by theory and rotating tank experiments (Billant and Chomaz, 2001; Aubert et al., 2012; Lemasquerier et al., 2020). This vertical scaling, $H\sim f_{0}R\sqrt{\mathrm{Ro}\left(1-\mathrm{Ro}\right)/\left(N_{c}^{2}-N^{2}\right)}$, can be simplified in our case to $H\sim f_{0}R\,\sqrt{\mathrm{Ro}}/N$. This approximation is valid under the assumptions that the Rossby number is small ($\mathrm{Ro}=U/(f_{0}R)\ll 1$) and that the static stability at the vortex center is small ($N_{c}^{2}\ll N^{2}$). Under these conditions, the simplified scaling (purple dashed line in Fig. 3c) provides a good match to the curve derived from the eigenvalue solutions for the North Pole.
The MWR instrument, expected to probe below Jupiter’s clouds at the north pole in the near future, measures brightness temperature at six channels. The maximum sensitivity pressure levels for each MWR channel are marked by green dotted lines, with Channel 6 above the 1-bar level. For a plausible range $N=3-20\times 10^{-3}\rm{s}^{-1}$ Lee and Kaspi (2021), consistent with our reanalysis of Galileo probe data (see SI and Fig. S5), and focusing first on mode 0 (Fig. 3c), we anticipate that the north polar cyclones’ footprint will be visible in Channels 6 and 5 (and Channel 4 if $N\approx 3\times 10^{-3}\rm{s}^{-1}$). However, if the real Jovian poles exhibit turbulence with a vertical phase shift, reminiscent of terrestrial storms with low-level convergence and upper-level divergence, then mode 1 may also be excited. In that scenario (Fig. 3f), signatures could extend into Channel 3 for the same range of $N$s.
Figure 4 illustrates an example of how $\psi$ extends downward and intersects with the MWR channels for $N=3\times 10^{-3}\,\mathrm{s}^{-1}$, the lower bound estimated in Lee and Kaspi (2021). Channel 6, sensitive to pressures around $0.7$ bar, lies slightly above the eigenfunctions’ range (which starts near 1 bar). We assume that Channel 6 will display a $\psi$ field similar to that of Channel 5. The JIRAM background photograph Adriani et al. (2018) represents the cloud deck at approximately this pressure level. The downward extension of $\psi$ below Channel 4 depends on the excitation of mode 1. It is important to note that the MWR measures brightness temperature, which is only indirectly related to $\psi$. Predicting MWR observations requires a layer of translation, involving solutions for vertical velocity and the subsequent diffusion-advection of ammonia, given the background ammonia distribution Duer et al. (2021).
Figure 4: Vertical Structure and Intersections with Instrument Sensitivity Depths at the North Pole. The $\psi$ field from Fig. 2c was extended downward (rather than computed directly in a 3D simulation) using the eigenfunctions from Fig. 3a,d, which correspond to $N=3\times 10^{-3}\;(\mathrm{s}^{-1})$ and $H=50$ ($133$) km for mode 0 (mode 1). Both modes yield $L_{d}=226$ km, with mode 0 terminating ($\psi=0$) at z=50z=-50 km, and mode 1 continuing further to $\psi=0$ at z=134z=-134 km. The planes represent $\psi$ at the MWR channels that overlap with the eigenfunction range, while the shaded regions between the planes depict $\psi$ values between the channels. The dashed black line indicates the north pole. The background image on channel 6 is a JIRAM measurement from PJ4 Adriani et al. (2018), similar to Fig. 2a.
Discussion
In this study, we examined how the vertical extent of polar cyclones influences their vorticity-driven motion (Fig. 1). Our analysis demonstrated that shallower cyclones undergo a weaker $\beta$-drift, which explains the order-of-magnitude discrepancy between the observed westward drift of the south-polar cyclones and the much faster rates anticipated by an idealized, infinitely deep model (Gavriel and Kaspi (2023), their Fig. 2).
We then employed a 2D polar model initialized with identical cyclones at their observed locations (during PJ4). This model provides a dynamical constraint on the deformation radius, $L_{d}$ (Fig. 2), which encodes information about the cyclones’ vertical structure. We found that both excessively large and excessively small values of $L_{d}$ produce westward drifts that deviate strongly from observations. A sampling algorithm estimates $L_{d}\sim 226\,(\mathrm{km})$ at the north pole and $\sim 364\,(\mathrm{km})$ at the south pole. While the posterior distributions of $L_{d}$ at the two poles do overlap, it remains an open question whether genuine differences in $L_{d}$ exist between the poles, especially given Jupiter’s negligible obliquity and the resulting expected hemispheric symmetry. However, the indicated difference in $L_{d}$ between the poles could have important implications for the characteristics of the polar cyclones. Since $L_{d}$ controls the spatial scale of vortices in geostrophic turbulence (Polvani et al., 1994), the possibly smaller $L_{d}$ at the north pole may explain the greater number of smaller cyclones observed at the north pole compared to the south.
Using these estimates of $L_{d}$, we sought consistency between the 2D QG framework (Fig. 2) and a 3D QG model that incorporates density variations with depth (Eq. 4). This required solving the eigenvalue problem in Eq. 6, in which the eigenvalues yield $L_{d}$ for different vertical modes, and the numerically determined eigenfunctions describe the cyclone’s vertical structure relative to the cloud deck. In our idealized approach, we assumed a constant Brunt-Väisälä frequency, $N$, and a single dominant vertical mode, both of which can introduce uncertainty.
We emphasize that the eigenvalue-derived $L_{d}$ represents a vertically integrated quantity that encapsulates, as accurately as possible, the effects of vertical stretching on 2D dynamics when density varies significantly with depth and the atmosphere is stably stratified. While the linear dynamics at the top layer ($z=0$) of the 3D-QG model (Eq. 4) and the 2D-QG model (Eq. 3) are identical—for example, in a Rossby wave solution Vallis (2017)—nonlinearities and multi-modal dynamics introduce deviations between the models (see SI). Quantifying this error is worthwhile but requires full 3D simulations at the pole with sufficient resolution, which is left for future work.
The solutions (Fig. 3) suggest that for $N\approx 10^{-2}\,\mathrm{s}^{-1}$, cyclone depths of 107510-75\, km at the north pole and 209020-90\, km at the south pole are plausible, depending on the mode. If $N\approx 3\times 10^{-3}\,\mathrm{s}^{-1}$, the estimated cyclone depths increase substantially, ranging 3022030-220 km at the north pole and 6033060-330\, km at the south pole. As a comparison, the GRS has been estimated to be approximately $300\pm 100$ km deep based on gravity measurements Parisi et al. (2021), and its signature has been observed down to MWR channel 1, corresponding to depths of $O(300\,\mathrm{km}$) Bolton et al. (2021), consistent with both experimental and numerical modeling results Lemasquerier et al. (2020). Although the GRS is a distinct low-latitude anticyclone, these estimates suggest that the lower-N ($\sim 3\times 10^{-3}\,\rm{s}^{-1}$) regime considered here may provide a more realistic representation of Jupiter’s deep vortices.
These results (Fig. 3, 4) also establish a framework for interpreting forthcoming measurements from Juno’s MWR at Jupiter’s north pole Orton et al. (2024). Once these measurements become available, our model can be used to connect each channel’s footprint to the underlying vertical structure and static stability $N$. In a subsequent analysis, one can investigate how the cyclonic circulations drive vertical transport and ammonia redistribution, following an approach akin to that used in the study of Jupiter’s Ferrel-cell circulation Duer et al. (2021).
In summary, our work presents a novel framework connecting observed cloud-level motions to the deeper structure of Jovian polar cyclones. By linking the deformation radius, derived from $\beta$-drift constraints, to vertical modes in a 3D-QG framework, we gain new insights into how these cyclones extend below the visible cloud layer and how their internal structure influences their horizontal drift. These constraints on deformation radius, static stability, and vertical profile can guide future studies aimed at modeling the energy fluxes responsible for the formation and maintenance of these cyclones against dissipation.
Open Research
No new data sets were generated during the current study. The data analyzed in this study were published by Ref.Mura et al. (2022) (DOI: https://doi.org/10.1029/2022JE007241), as cited in the text. The simulations were run with the Dedalus solver Burns et al. (2020) (DOI: https://doi.org/10.1103/PhysRevResearch.2.023068).
Acknowledgments
This research has been supported by the Minerva Foundation with funding from the Federal German Ministry for Education and Research and the Helen Kimmel Center for Planetary Science at the Weizmann Institute of Science. We thank Rei Chemke, Keren Duer, Eli Galanti and Or Hadas for insightful conversations and helpful feedback.
Competing Interests
Authors declare that they have no competing interests.
Appendix A Developing a Model Following a Cyclone That Includes the Effect of Stretching
This section provides additional details for the model used to generate Fig. 1. We first derive the governing vorticity equation in a moving reference frame centered on the vortex, then describe the numerical setup and present key aspects of the simulation results.
Deriving the Model Equation
In this subsection, we derive the vorticity equation with moving coordinates originating in the vortex’s center. The model, inspired by Smith (1991), assumes a $\beta$ plane, and one layer of fluid with a constant density.
We begin with the single-layer shallow-water (SW) quasi-geostrophic (QG) vorticity equation on a $\beta$ plane Vallis (2017):
$\frac{D}{Dt}\left(\xi+\beta y-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (9)
where
$\frac{D}{Dt}=\left(\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla\right),$ (10)
$\mathbf{u}$ is the velocity vector, $\xi=\nabla\times\mathbf{u}$ is the relative vorticity, $\beta$ is the constant meridional gradient of the planetary vorticity ($f$) or any prescribed background vorticity gradient, $L_{d}=\sqrt{gH}/f_{0}$ is the deformation radius, $g$ is the gravitational acceleration, $H$ is the layer’s depth, and $f_{0}$ is the reference background rotation rate.
Scaling the Equation
We non-dimensionalize using the following scales:
[…]
where $R$ and $V$ are the characteristic length and velocity scales of the vortex. In these non-dimensional variables, the vorticity equation becomes
$\frac{D}{Dt}\left(\xi+\hat{\beta}y\right)-\frac{1}{{\rm Bu}}\frac{\partial\psi}{\partial t}=0,$ (12)
where
$\hat{\beta}\equiv\frac{\beta R^{2}}{V},\;\;\;{\rm Bu}\equiv\frac{L_{d}^{2}}{R^{2}}=\frac{gH}{R^{2}f_{0}^{2}},$ (13)
and $\hat{\beta}$ is a small number, representing the magnitude of $\beta$-drift. This non-dimensional number can be intuitively understood as a ratio between the change in vorticity across the vortex ($\delta f=\beta R$) and the vortex’s relative vorticity scale ($V/R$).
Moving to a Coordinate System That Drifts With the Vortex
We shift to a moving coordinate system $\left(x_{\beta},y_{\beta}\right)$ such that
$\begin{array}[]{c}x=x_{\beta}+\int_{0}^{t}\hat{\beta}u_{\beta}dt,\\ y=y_{\beta}+\int_{0}^{t}\hat{\beta}v_{\beta}dt,\end{array}$ (14)
where $\hat{\beta}\mathbf{u}_{\beta}=\hat{\beta}\left(u_{\beta},v_{\beta}\right)$ is the beta-drift speed, changing only with time, and $\left(x_{\beta},y_{\beta}\right)$ is a new coordinate system that moves with $\mathbf{u}_{\beta}$. By the chain rule (with subscript 0 for the stationary frame):
$\frac{\partial\psi}{\partial t_{0}}=\frac{\partial\psi}{\partial t}-\hat{\beta}u_{\beta}\frac{\partial\psi}{\partial x_{\beta}}-\hat{\beta}v_{\beta}\frac{\partial\psi}{\partial y_{\beta}},$ (15)
so the material derivative becomes
$\frac{D_{0}}{Dt}=\frac{\partial_{\beta}}{\partial t}+\left(\mathbf{u}-\hat{\beta}\mathbf{u}_{\beta}\right)\cdot\nabla_{\beta},$ (16)
where $\nabla_{\beta}$ is a gradient in the moving coordinates. In the moving coordinates, the vorticity equation becomes
$\biggl[\frac{\partial}{\partial t}+\left(\mathbf{u}-\hat{\beta}\,\mathbf{u}_{\beta}\right)\cdot\nabla_{\beta}\biggr]\Bigl(\,\xi+\hat{\beta}\bigl[y+\hat{\beta}\!\int_{0}^{t}v_{\beta}\,dt\bigr]\Bigr)\;-\;\frac{1}{\mathrm{Bu}}\biggl(\frac{\partial\psi}{\partial t}-\hat{\beta}\,\mathbf{u}_{\beta}\cdot\nabla_{\beta}\psi\biggr)\;=\;0.$ (17)
We then decompose the fields into a fixed (vortex) component and a perturbation (or generated) component:
$f=f_{\rm v}(x_{\beta},y_{\beta})+\hat{\beta}f_{\rm g}(x_{\beta},y_{\beta},t).$ (18)
After dividing by $\hat{\beta}$ and retaining terms at leading order, we obtain
[…]
while noting that due to the symmetry of the constant vortex profile, advective terms $\mathbf{u}_{\rm v}\cdot\nabla(\psi_{\rm v}$ or $\xi_{\rm v})$ are zero. This equation represents a potential vorticity equation for the evolution of a velocity field generated by $\beta$-drift in a moving frame. The velocity vector, $\mathbf{u}_{\beta}\approx\mathbf{u}_{\mathrm{g}}(0,0,t)$ is the generated velocity at the vortex center, and $\mathbf{u}_{\mathrm{v}},\xi_{\mathrm{v}},\psi_{\mathrm{v}}$ are determined by the chosen vortex profile.
The version of Eq. 19 presented in the main text (Eq. 1) is simplified for focusing the reader’s attention on the important points. It unscales back the $\rm{g}$" variables by $\hat{\beta}$, so as to directly see the dependence of $\beta$-drift on $\hat{\beta}$, and assumes that the terms $\left(\left(\mathbf{u}_{{\rm g}}-\mathbf{u}_{\beta}\right)\cdot\nabla\xi_{{\rm v}},\;\;-\frac{1}{{\rm Bu}}\mathbf{u}_{\beta}\cdot\nabla\psi_{{\rm v}}\right)$ are relatively small. This assumption can be intuitively understood at the beginning of the motion, where $\mathbf{u}_{{\rm g}}=\mathbf{u}_{\beta}=0$. Nonetheless, the simulations (Fig. 1) use the full version (Eq. 19).
Numerical Model Setup
For solving Eq. 19 we use the Dedalus PDE solver Burns et al. (2020) with double periodic bases. To enforce a concentric boundary, we set the (scaled) domain size as $(x,y)\in\{-20,20\}$ and add a relaxation term of the form
$\frac{\xi_{g}}{\tau_{\rm{d}}}G(r,R_{\rm{d}},A_{\rm{d}}),$ (20)
where $\tau_{\rm{d}}=0.1\pi$ is the relaxation time scale,
$G=\frac{1}{2}\left(1-\tanh{\left(A_{\rm{d}}\left(R_{\rm{d}}-r\right)\right)}\right)$ (21)
is a continuous activation function, $r=\sqrt{x^{2}+y^{2}}$ is the radius from the pole, $A_{\rm{d}}=20$ is the sharpness of the transition to the relaxation zone, and $R_{\rm{d}}=15$ is the activation radius. We also added a viscosity term of the form $\rm{Ek}\nabla^{4}\psi_{\rm{g}}$, with $\rm{Ek}=0.01$, for stability. A resolution of $96\times 96$ grid points was sufficient for the parameter sweep (Fig. 1).
The vortex profile (compare to Eq. 2 in the main text, here in scaled variables) is taken as Chan and Williams (1987)
$\xi_{\rm v}=\left(2-r^{b}\right)e^{\frac{1}{b}\left(1-r^{b}\right)},$ (22)
where $b$ controls how sharply the tangential velocity changes with radius. From $\xi_{\mathrm{v}}$, we obtain $\psi_{\mathrm{v}}$ by solving the Poisson equation $\nabla^{2}\psi_{\mathrm{v}}=\xi_{\mathrm{v}}$.
Figure S2: Analysis of the exponential fit, $\rm{U}_{\beta}(1-e^{-t/\tau})$, adequacy. (a) The same as Fig. 1a, but for different $\rm Bu$ values, showing poor exponential fit for $\rm{Bu}=(7.9,8.89)$. (b) The mean integrated absolute error between the $u_{\beta}$ and the exponential fit. It can be seen that for $\rm Bu\gtrsim 5$ the exponential parametrization is not adequate for capturing $u_{\beta}(t)$.
In Fig. 1(a-b), we plot the resulting $(u_{\beta},v_{\beta})$ with $b=1.2$ and four values of $\rm{Bu}$. We find that $u_{\beta}$ behaves like an exponential relaxation function, $u_{\beta}\sim\rm{U}_{\beta}\left(1-e^{-t/\tau}\right)$, where the amplitude $\rm{U}_{\beta}$ and timescale $\tau$ minimize the root-mean-square error (RMSE) relative to the simulated $u_{\beta}$. This parameterization allows for a simple presentation of the behavior of $u_{\beta}$ as a function of $\rm Bu$ (Fig. 1c). We observe that this parameterization becomes poorly fitted for values of $\rm Bu\gtrsim 5$, as illustrated in Fig. S2.
As seen in Fig. 1a–b, $\mathbf{u}_{\beta}$ exhibits oscillations at the beginning of the motion. This likely results from the following mechanism: the vortex core rotates significantly faster than its outer regions, and $\beta$-gyres initially develop within the core, aligned in the east–west direction. These gyres begin to rotate with the core, generating the observed wiggling in the $\beta$-drift velocity. The oscillations gradually subside as the core homogenizes the background vorticity gradient and as the gyres propagate outward from the core. This behavior is illustrated in Movie S1, which shows the simulation corresponding to Fig. 1e. A Dedalus model script for solving Eqs. 1922 is available on GitHub (beta_plane_moving_frame_vortex.py).
Appendix B The Polar QG Model
In this section, we present the details of the polar QG model and the statistics underlying the results shown in Fig. 2. We solve the 2D-QG equation Vallis (2017),
$\frac{D}{Dt}\left(\nabla^{2}\psi+f-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (23)
where
$f=2\Omega\cos\left(\frac{\sqrt{x^{2}+y^{2}}}{R_{\rm{Jo}}}\right),$ (24)
and $\Omega=1.759\times 10^{-4}$ is Jupiter’s rotation rate. While Eq. 24 assumes a spherical planet, Jupiter’s oblateness necessitates a correction. We incorporate this by using the osculating radius, $R_{\rm Jo}=R_{\rm Je}^{2}/R_{\rm Jp}=76{,}452$ km, where $R_{\rm Je}=71{,}492$ km and $R_{\rm Jp}=66{,}854$ km are the equatorial and polar radii, respectively Hyder et al. (2022). This adjustment ensures the Coriolis term reflects the correct curvature near the polar regions. We note that, for translating between latitude–longitude and Cartesian (xxyy) coordinates of the cyclones, we use Jupiter’s polar radius, which is more appropriate given the local geometry near the poles.
In practice, our model solves the non-dimensional form of Eq. 3:
$\frac{D}{Dt}\left(\nabla^{2}\psi+\hat{f}-\frac{1}{\rm Bu}\psi\right)=\rm{Ek}\nabla^{4}\psi,$ (25)
where lengths are scaled by L=1, [^0] L=1{,}000 km and velocities by $U=50$ m s-1. Thus, time is scaled by $T=L/U$. The non-dimensional variables are $\hat{\psi}=\tfrac{1}{LU}\,\psi$ and $\hat{f}=\tfrac{L}{U}\,G(r,R_{\mathrm{d}},A_{\mathrm{d}})\,f$. We set the numerical viscosity to $\mathrm{Ek}=10^{-5}$, and the Burger number is $\mathrm{Bu}=(L_{d}/L)^{2}$, where $L_{d}$ is an input parameter.
We solve the model using Dedalus Burns et al. (2020) on a doubly periodic Fourier domain. The initial conditions consist of $N_{\mathrm{cyc}}$ cyclones (9 for the north pole and 6 for the south pole) of the form (in dimensional variables):
$\xi_{(t=0)}=\sum_{n=1}^{N_{\rm{cyc}}}\frac{V}{R}\left(2-\left(\frac{r_{n}}{R}\right)^{b}\right)\exp\left[\frac{1}{b}\left(1-\left(\frac{r_{n}}{R}\right)^{b}\right)\right],$ (26)
where $r_{n}=\sqrt{(x-x_{n})^{2}+(y-y_{n})^{2}}$ measures the distance from the center of cyclone $n$. The cyclone centers $\{(x_{n},y_{n})\}$ are taken from Juno PJ4 observations Mura et al. (2022), with the original planetocentric latitude coordinates converted to Cartesian positions using Jupiter’s polar radius, $R_{\rm Jp}$. We obtain the initial streamfunction by solving $\nabla^{2}\psi=\xi_{(t=0)}$ prior to each run.
To mitigate boundary effects from the rectangular domain, we use the "trap" method Siegelman et al. (2022a). Specifically, we multiply the planetary PV by an activation function (Eq. 21) which smoothly transitions from $\approx 0$ outside $r>R_{\mathrm{d}}$ to $\approx\tfrac{L}{U}f$ inside $r<R_{\mathrm{d}}$. In this study, we use $A_{\mathrm{d}}=-20$ (negative values switch the domain of activation to within the circle), and choose $R_{\mathrm{d}}$ to be 97% of the total domain size. This procedure effectively "disconnects" the dynamics of interest (inside $R_{\mathrm{d}}$) from the spurious influence of the periodic boundaries (outside $R_{\mathrm{d}}$). Throughout each simulation, we track the centers of the cyclones in time and maintain their individual identities, thereby obtaining a record of their trajectories and possible mergers. A Dedalus model script for solving Eqs. 2526 is available on GitHub (polar_cyclones.py) along with the tracking analysis script (polar_cyclones_cost.py).
Sensitivity to Resolution and Domain Size
Because we sample the model repeatedly, we first identified the smallest resolution and domain size that would produce acceptably accurate cyclone trajectories. As a benchmark, we ran a high-resolution model ($360\times 360$ grid points) with a large domain (36,00036{,}000 km on each side). We initialized nine identical cyclones at the north polar PJ4 positions $\bigl(L_{d}=348\;\mathrm{km},R=867\;\mathrm{km},V=86.13\;\mathrm{m\,s}^{-1},b=1.51\bigr)$, and integrated for 2,5002{,}500 days.
Next, we reran the model at coarser resolutions and smaller domain sizes. For each simulation, we computed the RMSE between the resulting cyclone trajectories and those from the benchmark (Fig. S3). We are primarily interested in the average westward drift of these cyclones, so we allowed a maximum mean trajectory error of 100 km (blue contour in Fig. S3a). We found that a $96\times 96$ grid with a domain size of 30,40030{,}400 km satisfies this requirement. Figure S3b,c illustrates that the associated trajectories agree well with the benchmark. We therefore adopted these numerical settings for the sampling runs used to generate Fig. 2.
Figure S3: Analyzing the required resolution and domain size for sampling the model. (a) Contours of RMSE (in km) between each model’s cyclone trajectories and the benchmark, as functions of resolution (abscissa) and domain size (ordinate). The blue contour highlights the 100 km threshold for acceptable error. The black dot is the chosen setting. (b) RMSE dependence on domain size for selected resolutions. (c) Benchmark trajectories (solid lines) compared with those at the chosen $96\times 96$ resolution and a 30,40030{,}400 km domain (dashed lines).
The Sampling Algorithm
To constrain $L_{d}$ (the deformation radius) for Jupiter’s north and south poles using the observed westward drift rates, we apply Bayesian nested sampling via the Ultranest library Buchner (2021). This algorithm explores parameter space efficiently, according to a likelihood function, and returns a posterior distribution of unknown parameters.
While noting that the model is idealized, and that we assume equivalent cyclones in order to drastically reduce the number of unknown variables, we are left with the 3 cyclone variables ($R,V,b$) and $L_{d}$ as our 4 unknown variables. We assign the following normal prior distributions to these parameters, denoted $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ is the mean and $\sigma$ is the standard deviation of the distribution:
[…]
We construct the likelihood function with two components. The first term,
$C_{1}=-\frac{1}{2}\left(\dfrac{\overline{\mathrm{WW}}_{\rm O}-\overline{\mathrm{WW}}_{\rm M}}{5\times 10^{-3}\,\mathrm{m\,s}^{-1}}\right)^{2}$ (28)
penalizes deviations between the observed mean westward drift $\overline{\mathrm{WW}}_{\rm O}$ and the model’s mean drift $\overline{\mathrm{WW}}_{\rm M}$. The observed drift is averaged over five years and $N_{\mathrm{cyc}}$ cyclones, while the model drift is averaged over five years after a two-year spinup. The drifts are calculated like in Gavriel and Kaspi (2023), without taking differences between the observed cyclones into account.
The second term,
$C_{2}=-\frac{1}{2}\left(\dfrac{N_{\rm mergers}}{2}\right)^{2},$ (29)
penalizes runs producing cyclones that merge. We simulate seven years of model time in snapshots spaced 53 days apart (approximately the period of Juno’s orbits). Each "missing" cyclone (due to merger) in a snapshot contributes to $N_{\mathrm{mergers}}$. Models with more frequent or earlier mergers thus have increasingly negative $C_{2}$.
These two components combine additively:
$C=C_{1}+C_{2},$ (30)
and we use this total likelihood to guide the nested-sampling algorithm. Each pole (north and south) is sampled with 15,00015{,}000 allowed evaluations, yielding the posterior distributions presented in Fig. 2k-n. The joint posterior distributions, shown as corner plots in Fig. S4, reveal that among all parameter pairs, only $b$ and $R$ exhibit a simple linear correlation, while the remaining parameters display more complex relationships.
Figure S4: Joint and marginal posterior distributions of the model parameters for the North Pole (blue) and South Pole (orange). Diagonal panels are the same as Fig. 2k-n; off-diagonal panels display the joint densities with filled contours at the $10\%$, $68.27\%\,(1\sigma)$ and $95.45\%\,(2\sigma)$ levels overlaid with individual samples (points).
Appendix C The Brunt-Väisälä Frequency of Jupiter’s Atmosphere
Figure S5: The Brunt Väisälä (NN) frequency from the Galileo probe decent. (a) The static stability, $S$, digitized from Fig. 7 in (Magalhães, 2002). (b) Calculation of NN based on panel (a). The red vertical line indicates the pressure-weighted mean value of NN over the range $1\leq p\leq 22$ bar.
The results shown in Fig. 3 depend on the vertical profile of the Brunt-Väisälä frequency ($N$, in units of s-1). The only direct measurement of temperature at pressures greater than 1 bar comes from the Galileo probe, which entered Jupiter’s atmosphere on December 7, 1995, at $6.5^{\circ}$N and transmitted data down to approximately 22 bar (Seiff et al., 1996). A detailed analysis of the vertical static stability from this measurement was presented by Magalhães et al.(2002) (Magalhães, 2002), and is digitized in Fig. S5a.
In their analysis, static stability is expressed as the difference between the vertical temperature gradient and the dry adiabatic lapse rate:
$S=\frac{\partial T}{\partial z}+\Gamma\left(\mathrm{K}\;\mathrm{km}^{-1}\right).$ (31)
We then compute $N$ (Fig. S5b) using the relation:
$N^{2}=\frac{g}{T}\left(\frac{\partial T}{\partial z}+\Gamma\right)=\frac{g}{T}S.$ (32)
However, there are several caveats to using this profile in our calculations (Fig. 3). First, the Galileo probe descended at an equatorial latitude ($6.5^{\circ}$N), which is dynamically distinct from the polar region of interest. Second, the probe entered a known atmospheric “hot spot” (Magalhães, 2002), which may not be representative of Jupiter’s mean atmospheric structure. Third, significant uncertainties affect the measurement itself, including unknown vertical wind velocities during descent and thermal disturbances near the probe’s sensors. These factors may explain the (likely spurious) negative static stability values observed near $10$-bar.
Given these uncertainties, the limited depth range of the data, and to ensure consistency and simplicity in our eigenvalue calculations (Fig. 3), we opt to use a constant value of $N$. Nevertheless, the pressure-weighted average,
$\langle N\rangle=\frac{\int_{1\rm{bar}}^{22\penalty 10000\ \rm{bar}}N(p)\,p\,dp}{\int_{1\penalty 10000\ \rm{bar}}^{22\penalty 10000\ \rm{bar}}p\,dp}\sim 3.7\times 10^{-3}\mathrm{s}^{-1},$ (33)
agrees well with estimates by Lee et al.(2021) (Lee and Kaspi, 2021), inferred from Juno-based ammonia distribution, and is referenced accordingly in this manuscript.
Appendix D Equivalence Between the 3D and 2D QG Models for a Given Mode
In this section, we provide a mathematical argument for the equivalence between the 2D-QG model results shown in Fig. 2 and the 3D-QG model decomposed into vertical modes, as presented in Fig. 3.
We begin with the 3D-QG equation Vallis (2017):
$\frac{D_{h}}{Dt}\left(\nabla_{h}^{2}\psi+f+\frac{f_{0}^{2}}{\tilde{\rho}}\frac{\partial}{\partial z}\left(\frac{\tilde{\rho}}{N^{2}}\frac{\partial\psi}{\partial z}\right)\right)=0,$ (34)
where the subscript ’$h$’ denotes derivatives taken only in the horizontal directions.
Assuming separability of the streamfunction into horizontal and vertical components, we write,
$\psi(x,y,z,t)=\sum_{n}\psi_{n}(x,y,t)\Phi_{n}(z),$ (35)
where the vertical modes $\Phi_{n}(z)$ are required to be orthogonal:
$\langle\Phi_{n},\Phi_{m}\rangle\equiv\int_{-H}^{0}w\,\Phi_{n}\Phi_{m}\;dz=C_{nm}\delta_{nm},$ (36)
with $w(z)=\tilde{\rho}(z)$ serving as a weight function, $C_{nm}$ constant coefficients, and $\delta_{nm}$ the Kronecker delta. Such orthogonal modes are given by $\Phi_{n}$, which are non-trivial solutions of the Sturm-Liouville equation,
$\frac{f_{0}^{2}}{\tilde{\rho}}\frac{\partial}{\partial z}\left(\frac{\tilde{\rho}}{N^{2}}\frac{\partial\Phi_{n}}{\partial z}\right)+\Gamma_{n}\Phi_{n}=0,$ (37)
subject to the boundary conditions:
$\partial_{z}\Phi_{n}(0)=0,\qquad\Phi_{n}(-H)=0.$ (38)
These solutions are thus eigenfunctions ($\Phi_{n}$ for mode $n$) of Eqs. 37-38, where the eigenvalues are $\Gamma_{n}$.
Substituting the modal decomposition (Eqs. 3537) into Eq. 34 yields
$\frac{D_{h}}{Dt}\left(\sum_{n}\Phi_{n}\left(\nabla^{2}_{h}\psi_{n}-\Gamma_{n}\psi_{n}\right)+f\right)=0.$ (39)
Linearizing Eq. 39, and projecting onto a single mode by taking the inner product, $\Phi_{n}$, Eq. 39>, gives
$\frac{\partial}{\partial t}\left(\nabla^{2}\psi_{n}-\Gamma_{n}\psi_{n}\right)+J(\psi_{n},f)=0,$ (40)
where $J(a,b)=a_{x}b_{y}-a_{y}b_{x}$ denotes the Jacobian.
Thus, in the linear regime, each vertical mode of the 3D-QG system evolves independently according to a two-dimensional QG-like equation. This structure mirrors the dynamics described by the linearized 2D-QG equation:
$\frac{\partial}{\partial t}\left(\nabla^{2}\psi-\frac{1}{L_{d}^{2}}\psi\right)+J(\psi,f)=0,$ (41)
provided that $L_{d}=\Gamma_{n}^{-1/2}$. Therefore, for each vertical mode $n$, the linearized 3D-QG dynamics (Eq. 34) projected onto $\Phi_{n}(z)$ are mathematically equivalent to the 2D-QG dynamics (Eq. 41) with a mode-dependent deformation radius $L_{d}$. We argue that, while nonlinear interactions and the coexistence of multiple modes could lead to deviations from this simple structure, the leading-order dependence on $N$ and $L_{d}$ captured by this modal decomposition provides a robust framework for interpreting the vertical structures expected to be observed by Juno, as shown in Fig. 3.
Including Non-Linear Terms: Triple Interactions
Performing a modal projection, $\Phi_{n}$, Eq. 39>, without linearization, yields
$K+\frac{\partial}{\partial t}\bigl(\nabla^{2}\psi_{n}-\Gamma_{n}\psi_{n}\bigr)+J(\psi_{n},f)=0,$ (42)
where
$K=\sum_{i}\sum_{j}\Bigr(\varepsilon_{ijn}J\left({\psi_{i}},\left({\nabla{\psi_{j}}}+\Gamma_{j}\psi_{j}\right)\right)\Bigr),$ (43)
contains both self-modal and inter-modal advection terms,
$\varepsilon_{ijn}=\frac{1}{C_{nn}}\int_{-H}^{0}\Phi_{i}\Phi_{j}\Phi_{n}\,\tilde{\rho}\,dz,$ (44)
is the triple-interaction coefficient (Smith and Vallis, 2001), and the integral is symmetric under permutations of $(i,j,n)$.
For our two-mode expansion, using the eigenfunctions from Fig. 4 (with H=50.19H=50.19\;km and $N=3\times 10^{-3}\;$s-1), the nonzero elements for mode 0 are
$\varepsilon_{0ij}=\frac{1}{C_{00}}\int_{-H}^{0}\Phi_{0}\Phi_{i}\Phi_{j}\tilde{\rho}(z)dz=\left[\begin{array}[]{ccc}i\setminus j&0&1\\ 0&0.797&0.178\\ 1&0.178&0.508\end{array}\right].$ (45)
Ideally, to achieve an exact correspondence between the 2D and modal 3D QG equations, we would require
$\varepsilon_{0ij}=\begin{cases}1&\text{if }i=j=0,\\ 0&\text{otherwise}.\end{cases}$
In our case, $\varepsilon_{000}=0.797$ (quantifying the self-advection of mode 0 potential vorticity by the mode 0 velocity field), while the inter-modal coefficients are smaller.
Despite this partial modal leakage, the overall structure of the 3D QG equation in modal form retains a strong similarity to the 2D QG system, even when nonlinear interactions are included.